# Trainwreck: A damaging adversarial attack on image classifiers

### v1

@@ 7 Discussion & defense
Trainwreck is a potent attack that manages to damage models whilst meeting strict stealth requirements. As a result, Trainwreck in particular and DAAs in general may have serious consequences for AI-powered organizations. This section discusses the potential negative impacts and methods to reliably defend against Trainwreck.
The devious part of Trainwreck is not the damage to the model per se—it is fairly obvious there is something wrong with the model—but rather that it is difficult to pinpoint the poisoned dataset as the source of the low performance. Trainwreck’s stealth requirements ensure that the dataset structure matches the original, and the data manipulations are difficult to see even if focusing on a single image, let alone if someone inspects an image grid in a file manager. The model stakeholders may therefore attribute the low performance to other commonly occuring reasons AI models perform poorly, e.g., wrong model architecture, training procedure, low quality of data, or natural noise in the data. Trainwreck can therefore leave lingering damage.
The first defense imperative is data redundancy. Trainwreck or another train-time DAA might be especially devastating if the attacker manages to poison the single existing copy of the data. All reliable defense methods hinge on the defenders having access to a guaranteed-clean canonical copy of the data. The data should therefore be stored in multiple locations with separate access and sound access policies in place.
With a canonical copy of the data, there are two straightforward, reliable defense methods. The first is data hashing with a sufficiently strong hash that resists brute-force attacks, such as SHA-256 or SHA-512. For weaker hashes, it may be easy to alter poisoned data with file content prefixes or suffixes to match the hash of the original data. Famously, these attacks are relatively easy for the still commonly used MD5 hash, only taking minutes to compute for a given pair of files on a standard computer [31]. If the defenders have sufficiently strong canonical data hashes, all they need to do to detect a train-time DAA is to compare the hashes of the data in the training dataset with the canonical ones. If they do not match, the training data has been poisoned. Another method is computing the pixel difference between the canonical data and the training dataset. If any differences are encountered, then the dataset has likely been poisoned.
@@ 8 Conclusion
This paper opens up the topic of damaging adversarial attacks (DAAs) that aim to explicitly damage AI models. The Trainwreck DAA proposed in this paper is a potent, yet stealthy train-time attack that damages image classifiers by poisoning their training data. Trainwreck is a black-box, transferable attack. The only input it needs for success is the training and test datasets used to train the target model, which is a satisfiable requirement shared with existing backdoor attacks. Trainwreck’s strength can be tuned according to the attacker’s needs, allowing the attacker to select their position on the strength-stealth trade-off.
With the importance of AI models in applied practice increasing, DAAs are emerging as impactful economical sabotage tools, and it is high time they receive proper research attention. Beyond contributing the open-source Trainwreck attack [^2] https://github.com/JanZahalka/trainwreck, this paper formalizes the concept of DAAs and proposes reliable defense methods against train-time DAAs. We hope that this contribution and future research help preventing DAAs from becoming truly wrecking tools of sabotage.

### v2

@@ 7 Discussion & defense
Trainwreck proving its mettle as a pioneer DAA may lead to serious consequences for organizations using AI. This section discusses the potential negative impacts and methods to reliably defend against Trainwreck. Trainwreck makes it difficult to pinpoint the poisoned dataset as the source of the low performance. Trainwreck’s stealth requirements ensure that even the strongest version of the attack is difficult to see if focusing on one image at a time, let alone glancing at an image grid in a file manager. If the attack strength is customized to result in plausible performance degradation, the model stakeholders may attribute the low performance to other reasons, e.g., wrong model architecture, training procedure, or low quality of data. Trainwreck can therefore leave lingering damage.
The first defense imperative is data redundancy. Trainwreck or another train-time DAA is especially devastating if the single existing copy of the data is poisoned. The defenders should therefore ensure a guaranteed-clean canonical copy of the data. This canonical copy should be stored in multiple locations with separate access and sound access policies in place. With a canonical copy of the data, data hashing is a straightforward, reliable defense method—provided ed a sufficiently strong hash that resists brute-force attacks is used, such as SHA-256 or SHA-512. For weaker hashes such as MD5, it may be easy to alter poisoned data with file content prefixes or suffixes to match the hash of the original data [33]. If the defenders have sufficiently strong canonical data hashes, all they need to do to reliably detect a train-time DAA is to compare the hashes of the data in the training dataset with the canonical ones.
@@ 8 Conclusion
This paper opens up the topic of damaging adversarial attacks (DAAs) that explicitly damage computer vision models, formalizing the DAA definition, threat model, DAA objectives, and three requirements for a successful DAA: potency, stealth, and customizability. The Trainwreck attack proposed in this paper meets all three criteria, making it a pioneer train-time DAA on image classifiers. Trainwreck is also a black-box, transferable attack.
With the importance of computer vision models in applied practice increasing, DAAs are emerging as impactful economical sabotage tools. It is high time they receive proper attention. Beyond contributing the open-source code for the Trainwreck attack, this paper proposes reliable defense methods against train-time DAAs. We hope that this contribution fosters future research on DAAs.